Light of wavelength 586.5 nm illuminates a slit of width 0.78 mm.

(a) At what distance from the slit should a screen be placed if the first minimum in the diffraction pattern is to be 0.93 mm from the central maximum?
________ m

(b) Calculate the width of the central maximum.
________ mm
Answer #1

a) condition of minima = wavelength* D / a

so D = 0.93 * 10-3 * 0.78 * 10-3 / 586.5 * 10-9

D = 1.23 m

b) width of central maxima = 2*0.93= 1.86 mm